OpenStudy (anonymous):

find (f(x+h)-f(x))/h if f(x)=x^2+6x. I got 2x+6 please check my answer

myininaya (myininaya):

does h->0 ? \[\frac{(x+h)^2+6(x+h)-(x^2+6x)}{h}=\frac{x^2+2xh+h^2+6x+6h-x^2-6x}{h}\] \[=\frac{2xh+h^2+6h}{h}=2x+h+6\]

OpenStudy (anonymous):

no h doesn't approach 0 you are right Thanks!

myininaya (myininaya):

so you are in algebra?